# Majorana Fermions in spin up and down electronic complexes in spin-orbit coupled array of semiconductor quantum dots in proximity to $s$-type superconductor and in magnetic field

### v1

@@ III.1 Exact diagonalization in normal Fermion basis
We begin by introducing the ED method used to obtain the energy spectrum of the s-wave Hamiltonian. In ED, the Hilbert space is constructed in a configuration (occupation-number) basis. For an electronic system consisting of $N$ spinful orbitals, the total Fock space dimension is $2^{2N}$, since each orbital can be either empty or occupied for each spin projection. The basis states can therefore be written as
@@ III.2.1 Onsite Hamiltonian and its matrix form
$\chi_{1}=\mathscr{l}\begin{pmatrix}-\xi\\ 1\end{pmatrix},\qquad\chi_{2}=\mathscr{l}\begin{pmatrix}1\\ \xi\end{pmatrix}$, (15)
@@
$\xi=\frac{\Delta}{\mu+\sqrt{\mu^{2}+\Delta^{2}}},\qquad\mathscr{l}=\frac{1}{\sqrt{1+\xi^{2}}}.$ (16)
The matrix $A$ mixes particle and hole operators (via $\Delta$). To simplify the structure of the Hamiltonian, it is advantageous to work in a basis that diagonalizes the on-site BdG matrix. We therefore define new fermionic operators $d_{j\sigma}$ by expanding the original Fermions in the eigenvectors of $A$ as
$\begin{pmatrix}c_{j\uparrow}\\ c_{j\downarrow}^{\dagger}\end{pmatrix}=\chi_{1}d_{j\uparrow}+\chi_{2}d_{j\downarrow}^{\dagger}=\mathscr{l}\begin{pmatrix}-\xi d_{j\uparrow}+d_{j\downarrow}^{\dagger}\\ d_{j\uparrow}+\xi d_{j\downarrow}^{\dagger}\end{pmatrix}$. (17)
@@ III.2.2 Transformation of hopping and SOC terms
$\tilde{t}=-\frac{2t\mu\mathscr{l}^{2}\xi}{\Delta},\qquad\Delta_{t}=-2t\mathscr{l}^{2}\xi.$ (20)
@@
$\tilde{\alpha}=\frac{2\alpha\mu\mathscr{l}^{2}\xi}{\Delta},\qquad\Delta_{\alpha}=-2\alpha\mathscr{l}^{2}\xi.$ (22)
@@ IV Results and discussion
Figure 6 shows the Fermion parity splitting $\Delta E=|E_{0}^{even}-E_{0}^{odd}|$, where $E_{0}^{even}$ and $E_{0}^{odd}$ denote the ground-state energies in the even and odd parity sectors, respectively. The results are obtained from ED for $N=3$ and $6$, and are complemented by density-matrix renormalization group (DMRG) calculations extending the system size up to $N=40$, implemented using the ITensor library [49]. The parity splitting of the ground states $\Delta E$ decreases monotonically with system size, signaling a nonlocal finite-size splitting mechanism associated with Majorana hybridization. While ED at $N\leq 6$ can therefore give the impression of a crossover to a strongly hybridized regime, since $\Delta E$ remains sizable within the limited ED window. The DMRG results demonstrate that this is predominantly a finite-size effect, upon increasing $N$ well beyond the ED accessible sizes, $\Delta E$ continues to decrease.

### v2

@@ III.1 Exact diagonalization in normal Fermion basis
We begin by introducing the ED method used to obtain the energy spectrum of the s-wave Hamiltonian. In ED, the Hilbert space is constructed in a configuration (occupation-number) basis. For an electronic system consisting of NN spinful orbitals, the total Fock space dimension is $2^{2N}$, since each orbital can be either empty or occupied for each spin projection. The basis states can therefore be written as
@@ III.2.1 Onsite Hamiltonian and its matrix form
$\chi_{1}=l\begin{pmatrix}-\xi\\ 1\end{pmatrix},\qquad\chi_{2}=l\begin{pmatrix}1\\ \xi\end{pmatrix}$, (15)
@@
$\xi=\frac{\Delta}{\mu+\sqrt{\mu^{2}+\Delta^{2}}},\qquad l=\frac{1}{\sqrt{1+\xi^{2}}}.$ (16)
The matrix AA mixes particle and hole operators (via $\Delta$). To simplify the structure of the Hamiltonian, it is advantageous to work in a basis that diagonalizes the on-site BdG matrix. We therefore define new fermionic operators $d_{j\sigma}$ by expanding the original Fermions in the eigenvectors of AA as
$\begin{pmatrix}c_{j\uparrow}\\ c_{j\downarrow}^{\dagger}\end{pmatrix}=\chi_{1}d_{j\uparrow}+\chi_{2}d_{j\downarrow}^{\dagger}=l\begin{pmatrix}-\xi d_{j\uparrow}+d_{j\downarrow}^{\dagger}\\ d_{j\uparrow}+\xi d_{j\downarrow}^{\dagger}\end{pmatrix}$. (17)
@@ III.2.2 Transformation of hopping and SOC terms
$\tilde{t}=-\frac{2t\mu l^{2}\xi}{\Delta},\qquad\Delta_{t}=-2tl^{2}\xi.$ (20)
@@
$\tilde{\alpha}=\frac{2\alpha\mu l^{2}\xi}{\Delta},\qquad\Delta_{\alpha}=-2\alpha l^{2}\xi.$ (22)
@@ IV Results and discussion
Figure 6 shows the Fermion parity splitting $\Delta E=|E_{0}^{even}-E_{0}^{odd}|$, where $E_{0}^{even}$ and $E_{0}^{odd}$ denote the ground-state energies in the even and odd parity sectors, respectively. The results are obtained from ED for $N=3$ and 66, and are complemented by density-matrix renormalization group (DMRG) calculations extending the system size up to $N=40$, implemented using the ITensor library [49]. The parity splitting of the ground states $\Delta E$ decreases monotonically with system size, signaling a nonlocal finite-size splitting mechanism associated with Majorana hybridization. While ED at $N\leq 6$ can therefore give the impression of a crossover to a strongly hybridized regime, since $\Delta E$ remains sizable within the limited ED window. The DMRG results demonstrate that this is predominantly a finite-size effect, upon increasing NN well beyond the ED accessible sizes, $\Delta E$ continues to decrease.